OpenStudy (anonymous):

In the cylinder below, the radius, r, is 2 ft, and the altitude, h, is 10 ft. What is the volume of the cylinder?

OpenStudy (johnweldon1993):

So the volume of a cylinder can be found by using \[\large \pi r^2 h\] So for your question you have \[\large \pi (2)^2 \times 10 = \space ?\]